# Modeling and computing many-body electronic properties of twisted bilayer graphene with mechanical relaxation

### v1

@@ Abstract
We introduce and compute solutions of a many-body model of the electronic properties of twisted bilayer graphene which systematically accounts for the effects of structural relaxation. We model mechanical relaxation by coupling linear elasticity to a stacking energy that penalizes disregistry. Minimizers of the resulting functional are then input into a tight-binding model of twisted bilayer graphene, from which a single-particle continuum moiré-scale (Bistritzer-MacDonald-like) model is systematically derived. We then project this model together with a Coulomb electron-electron interaction term into the single-particle model’s flat moiré bands. We numerically compute Hartree-Fock ground states of this model, comparing the relative energies of competing many-body ground states.
@@ 3 Theory: Many-body Model
where we recall $\varepsilon_{m\mathbf{k}}$ are the bands of the Bistritzer-MacDonald Hamiltonian Eq. 15. The double counting subtraction $\hat{H}_{\text{sub}}$ has a similar form, also quadratic in the creation and annihilation operators; we present this in Section 3.2 after we have introduced the Hartree and exchange operators.
@@ 3.1 Hartree-Fock Theory
so that the 1-RDM is block-diagonal with respect to relative wave-vector $\mathbf{k}$, with $|\mathcal{N}|\times|\mathcal{N}|$ blocks. It is important to emphasize that condition (51) still allows for non-zero 1-RDM entries coupling wave-vectors $\mathbf{k}$ in valley $\mathbf{K}$ and $\mathbf{k}$ in valley $\mathbf{K}^{\prime}$ and vice versa. Recalling (21), condition (51) therefore allows for coupling between the time-reversal pair of total wave-vectors $\tilde{\mathbf{K}}+\mathbf{k}$ and $-(\tilde{\mathbf{K}}+\mathbf{k})$, so that the 1-RDM is not translation invariant in the sense of being block diagonal with respect to total wave-vector. States coupling wave-vectors in different valleys are known as intervalley coherent (IVC). Such states have been theoretically predicted as ground states outside of the chiral limit [6] and have been observed in experiments [36]. In the chiral limit it is known [46] that many-body ground states are always translation-invariant in the sense of (51). In particular, ground state 1-RDMs may have non-zero entries between relative wave-vectors $\mathbf{k}$ in each valley, but every other intervalley entry must be zero.
@@
$\hat{K}[P]$ $=\frac{1}{|\Gamma_{m}^{\mathcal{K}}|}\sum_{\mathbf{k},\mathbf{k}^{\prime}\in\mathcal{K}}\sum_{\mathbf{G}\in\mathcal{R}_{\mathrm{m}}^{*}}\sum_{m,n,m^{\prime},n^{\prime}\in\mathcal{N}}\hat{V}(\mathbf{k}^{\prime}-\mathbf{k}+\mathbf{G})[\rho_{\mathbf{k},\mathbf{k}^{\prime}}(\mathbf{G})]_{mn}[P(\mathbf{k}^{\prime})]_{nm^{\prime}}[\rho_{\mathbf{k}^{\prime},\mathbf{k}}(-\mathbf{G})]_{n^{\prime}m^{\prime}}\hat{f}_{m\mathbf{k}}^{\dagger}\hat{f}_{n^{\prime}\mathbf{k}},$ (53)
and the electron-electron contribution to the energy can be written:
@@ 3.2 The Subtraction Hamiltonian
where $P_{0}$ is a fixed 1-RDM. Previous work has shown that the choice of subtraction can modify the nature of many-body ground states [22]. In this work, we choose the “average” subtraction, which corresponds to setting $P_{0}\equiv\frac{1}{2}I_{4\times 4}$ for all $\mathbf{k}\in\mathcal{K}$.
@@ 3.3 Symmetry Order Parameter
where the additional minus sign is due to the fact that $\nu_{x}\mathcal{T}$ satisfies $(\nu_{x}\mathcal{T})H(\mathbf{k})=H(-\mathbf{k})(\nu_{x}\mathcal{T})$. The symmetry order parameter for $\nu_{y}\mathcal{T}$ has an identical formula to $\nu_{y}\mathcal{T}$ with the sewing matrix $B_{\mathbf{k}}(\nu_{x}\mathcal{T})$ replaced with $B_{\mathbf{k}}(\nu_{y}\mathcal{T})$.
@@ A.1 Intralayer terms
with a relaxation dependent hopping function

### v2

# Interacting Twisted Bilayer Graphene with Systematic Modeling of Structural Relaxation

@@ Abstract
Twisted bilayer graphene (TBG) has drawn significant interest due to recent experiments which show that TBG can exhibit strongly correlated behavior such as the superconducting and correlated insulator phases. Much of the theoretical work on TBG has been based on analysis of the Bistritzer-MacDonald model which includes a phenomenological parameter to account for lattice relaxation. In this work, we use a newly developed continuum model which systematically accounts for the effects of structural relaxation. In particular, we model structural relaxation by coupling linear elasticity to a stacking energy that penalizes disregistry. We compare the impact of the two relaxation models on the corresponding many-body model by defining an interacting model projected to the flat bands. We perform tests at charge neutrality at both the Hartree-Fock and Coupled Cluster Singles and Doubles (CCSD) level of theory and find the systematic relaxation model gives quantitative differences from the simplified relaxation model.
@@ 3 Theory: Many-body Model
where we recall $\varepsilon_{m\mathbf{k}}$ are the bands of the Bistritzer-MacDonald Hamiltonian Eq. 15. The double counting subtraction $\hat{H}_{\text{sub}}$ has a similar form, also quadratic in the creation and annihilation operators; we present this in Section 3.2 after we have introduced the Coulomb and exchange operators.
@@ 3.1 Hartree-Fock Theory
so that the 1-RDM is block-diagonal with respect to relative wave-vector $\mathbf{k}$, with $|\mathcal{N}|\times|\mathcal{N}|$ blocks.
It is important to emphasize that condition (51) still allows for non-zero 1-RDM entries coupling wave-vectors $\mathbf{k}$ in valley $\mathbf{K}$ and $\mathbf{k}$ in valley $\mathbf{K}^{\prime}$ and vice versa. Recalling (21), condition (51) therefore allows for coupling between the time-reversal pair of total wave-vectors $\tilde{\mathbf{K}}+\mathbf{k}$ and $-(\tilde{\mathbf{K}}+\mathbf{k})$, so that the 1-RDM is not block-diagonal with respect to total wave-vector. It can be checked directly that Hartree-Fock states corresponding to 1-RDMs which are block-diagonal with respect to total wave-vector, see (61)-(62) for example, are indeed invariant (up to phase) under translations in each co-ordinate by moiré lattice vectors. Hartree-Fock states corresponding to 1-RDMs which are not block-diagonal with respect to total wave-vector, see (63)-(64) for example, need not be. Intervalley coherent states such as (63)-(64) also display a factor of $\sqrt{3}$ enlargement of the unit cell at the atomic scale because of the interference of wavenumbers near to the distinct Dirac points [6].
States coupling wave-vectors in different valleys while satisfying (51) are known as intervalley coherent (IVC). IVC states have been theoretically predicted as ground states outside of the chiral limit [6] and have been observed in experiments [36]. In the chiral limit it is known [46] that many-body ground states are always translation-invariant in the sense of (51). In particular, ground state 1-RDMs may have non-zero entries between relative wave-vectors $\mathbf{k}$ in each valley, but every other intervalley entry must be zero.
@@
$\hat{K}[P]$ $=\frac{1}{|\Gamma_{m}^{\mathcal{K}}|}\sum_{\mathbf{k},\mathbf{k}^{\prime}\in\mathcal{K}}\sum_{\mathbf{G}\in\mathcal{R}_{\mathrm{m}}^{*}}\sum_{m,n,m^{\prime},n^{\prime}\in\mathcal{N}}\hat{V}(\mathbf{k}^{\prime}-\mathbf{k}+\mathbf{G})[\rho_{\mathbf{k},\mathbf{k}^{\prime}}(\mathbf{G})]_{mn}[P(\mathbf{k}^{\prime})]_{nm^{\prime}}[\rho_{\mathbf{k}^{\prime},\mathbf{k}}(-\mathbf{G})]_{n^{\prime}m^{\prime}}\hat{f}_{m\mathbf{k}}^{\dagger}\hat{f}_{n^{\prime}\mathbf{k}}$ (53)
with corresponding matrix representations $J[P]$ and $K[P]$, and the electron-electron contribution to the energy can be written:
@@ 3.2 The Subtraction Hamiltonian
where $P_{0}$ is a fixed 1-RDM. Previous work has shown that the choice of subtraction can modify the nature of many-body ground states [22]. In this work, we choose the “average” subtraction, which corresponds to setting $P_{0}\equiv\frac{1}{2}I_{4\times 4}$ for all $\mathbf{k}\in\mathcal{K}$. This choice of subtraction gives that the interacting Hamiltonian, $\hat{H}$, is positive semidefinite for the chiral model after an appropriate constant shift [45].
@@ 3.3 Symmetry Order Parameter
where the additional minus sign is due to the fact that $\nu_{x}\mathcal{T}$ satisfies $(\nu_{x}\mathcal{T})H(\mathbf{k})=H(-\mathbf{k})(\nu_{x}\mathcal{T})$. The symmetry order parameter for $\nu_{y}\mathcal{T}$ has an identical formula to $\nu_{x}\mathcal{T}$ with the sewing matrix $B_{\mathbf{k}}(\nu_{x}\mathcal{T})$ replaced with $B_{\mathbf{k}}(\nu_{y}\mathcal{T})$.
@@ A.1 Intralayer terms
with a disregistry dependent hopping function
@@ A.2 Interlayer terms
A.3 Energy scales and comparison with previous models
[…]
First Order Second Order Third Order & higher
Description Term $E$ (meV) Description Term $E$ (meV) Description Term $E$ (meV)
Graphene dispersion Dirac $v\Delta\mathbf{K}$ 176 Quad. dispersion AA $v_{1}\Delta\mathbf{K}^{2}$ 1
Quad. dispersion AB $v_{2}\Delta\mathbf{K}^{2}$ 4
Intralayer hops Shell 1 AB $\left|[A_{1}^{(\ell)}]_{\mathrm{A}\mathrm{B}}\right|$ 18 Shell 2 AB $\left|[A_{7}^{(\ell)}]_{\mathrm{A}\mathrm{B}}\right|$ 0.32
Shell 1 $k$-dep $\left|A_{1,\nabla}^{(\ell)}\right|\Delta\mathbf{K}$ 1 Shell 2 $k$-dep $\left|A_{7,\nabla}^{(\ell)}\right|\Delta\mathbf{K}$ 0.02
Shell 1 & 2 AA $\left|[A_{j}^{(\ell)}]_{\mathrm{A}\mathrm{A}}\right|$ 0.02
Interlayer hops Shell 1 AA $\left|[\tilde{T}_{1}]_{\mathrm{A}\mathrm{A}}\right|$ 79 Shell 1 k-dep $\left|\tilde{T}_{1,\nabla}\right|\Delta\mathbf{K}$ 3 Shell 2 $k$-dep $\left|\tilde{T}_{4,\nabla}\right|\Delta\mathbf{K}$ 0.14
Shell 1 AB $\left|[\tilde{T}_{1}]_{\mathrm{A}\mathrm{B}}\right|$ 113 Shell 2 $\left|\tilde{T}_{4}\right|$ 3 Shell 3 $k$-dep $\left|\tilde{T}_{7,\nabla}\right|\Delta\mathbf{K}$ 0.35
Shell 3 $\left|\tilde{T}_{7}\right|$ 12
Table 5: Approximate energy scales of the individual terms in the relaxed BM Hamiltonian. The indices correspond to momentum hops in the respective shells (see Fig. 1(C).) Terms in bold are included in the original BM model [5]. Terms in bold and italic are corrections included in the higher-order BM model [40].